Ironlove Inc. considers a project to supply a car manufacturer with 105,000 tons of steel plates annually for five years. It requires an initial investment of 42 million dollars in steel plate production machinery. The variable cost of the production is $430 per ton. The overhead cost will be 15 million dollars a year, and the company estimates the price of the steel plates at $950 per ton. To run this project, Ironlove needs to invest 8 million dollars in the net working capital at time zero.

The machinery belongs to Class 6, with a CCA rate of 15%. The company estimates that it can sell the machinery at a salvage value of 12 million dollars. Assume the company requires a 15% rate of return, and the tax rate is 40%.

a) What are the estimated OCF and NPV for this project? Should you pursue this project?

b) Suppose you believe that the accounting department’s initial cost and salvage value projections are accurate only to within ±20%; the marketing department’s steel price estimate is accurate only within ±15%, and the net working capital estimate is accurate only to within ±10%. What is your worst-case and best-case scenario for this project?

c) Suppose you are not confident about the car manufacturer’s actual steel plates requirement. What is the sensitivity of the project OCF to changes in the quantity supplied?

d) What about the sensitivity of NPV to changes in quantity supplied? Given the sensitivity number you calculated, what is the minimum level of output below which you wouldn’t want to operate?
Answer #1
(a)
Year Operating Cash Flow Cash outflow Net Cashflow Depreciation Net profit Tax Cashflow after tax PVF @
15 %
Discounted Value  
(a) (b) (c ) (d=b-c) (e ) (f=d-e) (g=f x 40%) (h=d-g) (i) (J=h x i)
                                      -                              -           5,00,00,000         -5,00,00,000                          -           -5,00,00,000                            -           -5,00,00,000          1.00         -5,00,00,000
                                       1         3,96,00,000                            -             3,96,00,000           31,50,000          3,64,50,000         1,45,80,000          2,50,20,000          0.87          2,17,56,522
                                       2         3,96,00,000                            -             3,96,00,000           58,27,500          3,37,72,500         1,35,09,000          2,60,91,000          0.76          1,97,28,544
                                       3         3,96,00,000                            -             3,96,00,000           49,53,375          3,46,46,625         1,38,58,650          2,57,41,350          0.66          1,69,25,355
                                       4         3,96,00,000                            -             3,96,00,000           42,10,369          3,53,89,631         1,41,55,853          2,54,44,148          0.57          1,45,47,774
                                       5         3,96,00,000                            -             3,96,00,000           35,78,813          3,60,21,187         1,44,08,475          2,51,91,525          0.50          1,25,24,640
                                       5         2,00,00,000                            -             2,00,00,000                          -            2,00,00,000            80,00,000          1,20,00,000          0.50              59,66,121
Net Present Value          4,14,48,957
Working Notes:
(i)Net cash inflow (Operating Cash flow):
Year Sales VC Fixed Overheads Net cash inflow
(+/(-) + - -
1-5         9,97,50,000         4,51,50,000           1,50,00,000       3,96,00,000
5 Salvage vale and WC recovery ( 12 + 8) = $ 2,00,00,000
Sales (105000 x 950 )         9,97,50,000
VC (105000 x 430 )         4,51,50,000
(ii) Initial cost Machine '(42*1000000) + WC (8*1000000)       5,00,00,000
(iii) Depreciation
Year Beg. Value Depreciation End. Value
(a) (b) (c=b x 15%) (d=b-c)
1         4,20,00,000             31,50,000           3,88,50,000
2         3,88,50,000             58,27,500           3,30,22,500
3         3,30,22,500             49,53,375           2,80,69,125
4         2,80,69,125             42,10,369           2,38,58,756
5         2,38,58,756             35,78,813           2,02,79,943